OpenStudy (cggurumanjunath):

integrate 3xsecxtanx

OpenStudy (anonymous):

\[\int\limits 3x \cdot \tan(x) \sec(x) . dx\] Integrated. :P

OpenStudy (cggurumanjunath):

?

OpenStudy (anonymous):

I think By parts, will work here..

OpenStudy (anonymous):

Take algebraic as First function or \(u\) or the other as \(v\).

OpenStudy (cggurumanjunath):

there r 3 functions !

OpenStudy (anonymous):

Oh, thanks for telling that.. :P

OpenStudy (cggurumanjunath):

:)

OpenStudy (usukidoll):

isn't antiderivative of secxtanx secx?

OpenStudy (anonymous):

Let us start without knowing where actually we will reach after this.. \[\implies (3x) \int\limits(\sec(x).\tan(x)) dx - \int\limits(3) \cdot \int\limits(\sec(x)\tan(x))dx \\ \]

OpenStudy (cggurumanjunath):

yes

OpenStudy (anonymous):

Yes I am just applying that only here..

OpenStudy (anonymous):

Refer to the attachment from WolframAlpha.

OpenStudy (kainui):

My shot at it: \[\Large \int\limits x \sec x \tan x dx = \int\limits x (\sec x)' dx \\ \Large x = u \ \ \ \ \ (\sec x)' dx = dv \\ \Large dx = du \ \ \ \ \ \sec x = v \\ \Large x \sec x - \int\limits \sec x dx \] Now how do you handle the next one? \[\Large \int\limits \sec x dx = \int\limits \sec x \left( \frac{\sec x + \tan x}{\sec x + \tan x} \right )dx \\ \Large \int\limits \frac{\sec^2x + \sec x \tan x }{\sec x + \tan x} dx \\ \Large \int\limits \frac{(\tan x)' + (\sec x) ' }{\tan x+ \sec x} dx \\ \Large \ln | \tan x + \sec x | +C\]

OpenStudy (anonymous):

I took the wrong path in between.. :(

OpenStudy (anonymous):

\[3x \sec(x) - 3 \int\limits (\sec(x)) \cdot dx\]

OpenStudy (cggurumanjunath):



OpenStudy (cggurumanjunath):